OpenStudy (eric_d):

Differentiate following with respect to x. a) e^1/3x

OpenStudy (eric_d):

@zzr0ck3r

OpenStudy (anonymous):

\[\large \frac{ d }{ dx } e^{\frac{ x }{ 3 }}?\]

OpenStudy (eric_d):

I don't know..I'm new to this..

OpenStudy (zzr0ck3r):

hes asking you to clarify what function you mean to differentiate.

OpenStudy (zzr0ck3r):

its hard to tell without parentheses

OpenStudy (eric_d):

ok

OpenStudy (anonymous):

Well is your expression \[\huge e^\frac{ 1 }{ 3x }~~~~\text{or}~~~e^ \frac{ x }{ 3 }\]

OpenStudy (zzr0ck3r):

or \(e^\frac{1}{3}x\)

OpenStudy (eric_d):

LHS

OpenStudy (eric_d):



OpenStudy (zzr0ck3r):

@iambatman has got you, I got to go eat ;)

OpenStudy (anonymous):

Ah ok cool so when you differentiate e^x you get the result of e^x, kind of crazy right? Since I am assuming you're sort of new to the notation of derivatives I will use "prime" meaning differentiate and it denoted as `'` \[\huge (e^{\frac{ 1 }{ 3 }x})'\] that little apostrophe represents we have to get the derivative. So we can use the "chain rule" here assuming you've learned this already, so the process would go sort of like this: \[\huge (e^{\frac{ 1 }{ 3 }x})' \implies \frac{ 1 }{ 3 }e^{\frac{ x }{ 3 }} \]

OpenStudy (anonymous):

iambatman after you are done with this question could you help me??

OpenStudy (anonymous):

I guess to show the whole process with the "fancy" notation you would let u=x/3 \[\frac{ d }{ dx }(e^{x/3}) = \frac{ de^u }{ du }\frac{ du }{ dx }\] where u = x/3, d/du e^u = e^u as I mentioned the derivative of e^x = e^x, so \[e^{x/3}\left( \frac{ d }{ dx }\left( \frac{ x }{ 3 } \right) \right) \implies \frac{ 1 }{ 3 }\left( \frac{ d }{ dx }x \right)e^{x/3} \implies \frac{ 1 }{ 3 }e^{x/3}\] since the 1/3 is a constant just factor it out. and derivative of x is just 1.

OpenStudy (anonymous):

If you don't understand a certain part just ask, I can explain it, it can look messy I guess.

OpenStudy (eric_d):

Ok

OpenStudy (eric_d):

Understood. Thanks @iambatman

OpenStudy (anonymous):

Np, glad you could understand!